OpenStudy (anonymous):

Given the function f(x) = 4(x+3) − 5, solve for the inverse function when x = 3.

jimthompson5910 (jim_thompson5910):

Replace f(x) with y. Then swap x and y. f(x) = 4(x+3) - 5 y = 4(x+3) - 5 x = 4(y+3) - 5 The next step is to solve for y. Are you able to solve for y?